. A girl delivering newspapers covers her route by traveling 3.00



blocks west, 4.00 blocks north, and then 6.00 blocks east. (a)



What is her resultant displacement? (b) What is the total



distance she travels?

1
Expert's answer
2021-11-08T16:36:20-0500


According to diagram

Total displacement


"AE=\sqrt{AD^2+DE^2}=\sqrt{4^2+3^2}=5 block"

Total distance

d=AB+BC+CE

d=3+4+6=13 block